OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = (8/x^2) + 4 How can I split it?

OpenStudy (zzr0ck3r):

f(x) = x + 4 g(x) = 8/x^2

OpenStudy (zzr0ck3r):

so f(8/x^2) = (8/x^2) + 4

OpenStudy (anonymous):

ok thank you

OpenStudy (zzr0ck3r):

you could say infinite answers f(x) = x and g(x) = (8/x^2) + 4 or f(x) = x+-2 g(x) = (8/x^2) + 6